OpenStudy (anonymous):

x + x = 2x what are some numbers i can plug in for this to be true

OpenStudy (amistre64):

x= 125844358635 comes to mind

OpenStudy (anonymous):

x + x = x + x x = x anything baby

OpenStudy (anonymous):

Any numbers..

OpenStudy (anonymous):

x • x = 2x how bout this one?

OpenStudy (anonymous):

and then x^2=-9

OpenStudy (anonymous):

and r they always never or sometime

OpenStudy (anonymous):

x^2 = 2x just 2 or 0. x^2 = -9 x^2 + 9 = 0 if you want complex numbers... (x+3i)(x-3i) x = +/- 3i

OpenStudy (anonymous):

\[2 \times 4 + 2 \times 4 = 2 \times 8.\]

OpenStudy (anonymous):

which one is that for?